# RoboMP: A Robotic Multimodal Perception-Planning Framework with Multimodal Large Language Models

### v1

@@ 4.5 Ablation Study
Effects of the Coarse-to-fine Retriever. We conduct ablation experiments to valid the performance of RoboMP with different settings of the retriever-augmented module. In particular, we compare 4 different implementations: (1) w/o CR removes the coarse retriever, directly using the rewriting module to order all programs in the codebase; (2) w/o IR removes the rewriting module, remaining the coarse retriever and semantic reorder; (3) w/o SR removes the reorder module, unsorting the retrieved demonstrations; (4) w/o RAMP removes the whole retriever, and takes programs from the codebase randomly as prompt examples. As depicted in Table 4, without either of these components, the average performance drops significantly. The reason is that when the context is composed of irrelevant examples, it is difficult for the planner to generate execution plans successfully.
@@
| RoboMP | ✓ | ✓ | ✓ | 89.0 | 85.9 | 86.8 | 68.0 | 82.4 |
| RoboMP | ✗ | ✓ | ✓ | 14.4 | 16.0 | 19.6 | 8.0 | 14.5 |
| RoboMP | ✓ | ✗ | ✓ | 87.7 | 86.1 | 86.7 | 49.0 | 77.4 |
| RoboMP | ✓ | ✓ | ✗ | 81.9 | 80.6 | 82.7 | 65.0 | 77.5 |
| RoboMP | ✗ | ✗ | ✗ | 10.5 | 9.5 | 11.6 | 3.0 | 8.7 |
@@
To investigate the robustness of RoboMP, we construct an extensive test set by enriching the original instructions using GPT4. Subsequently, we conduct experiments to evaluate the impact of this augmentation. As shown in Table 3, the results demonstrate the outstanding semantic understanding capability of our approach. In comparison to I2A, RoboMP is not confined to processing text instructions with fixed patterns. On the contrary, it possesses stronger generalization and robustness.
@@
Comparison of Multimodal Perceptors. To validate the capabilities of GCMP for intricate relationship references, we construct a perception test set. This set contains fundamental perception tasks that detect objects by names, and complex perception tasks that requires understanding complex referential expressions as illustrated in Section 3.1.1. Since the traditional visual perceptors used in the robotic manipulation, such as YOLO, cannot take the referential expressions to perceive objects, we compare our embodied perceptor GCMP with SAM+CLIP in I2A, and the general MLLM, i.e., Shikra (Chen et al., 2023). It can be seen from Table 7 that our GCMP outperforms the two models on both simple and complex referential perception by a large margin. Neither SAMP+CLIP or Shikra can work on the complex referential perception. The reason is that SAM+CLIP does not have a ability to understand the complex semantic reference and the general model Shikra struggles to generalize to complex embodied perception tasks.
@@ 4.6 Qualitative Results
We further present qualitative results in Figure 5, illustrating the planning and execution steps of RoboMP in comparison to other methods across three different kinds of tasks. It can be observed that only RoboMP successfully accomplishes all tasks. While all three methods perform well on VIMABench tasks, the end-to-end policy fails to generalize to real-world scenarios due to differences in the robot arm. Furthermore, SAM+CLIP schema is shown to unable to recognize referential objects with complex relationships. This leads I2A to only complete pick-and-place tasks for simple objects, resulting in failure in the second step of sim-to-real. In the generalization scenario, other methods are unable to complete the instruction. Only RoboMP, leveraging its robust language understanding and multimodal perception capabilities, successfully identify objects with intricate relationships. It adaptively generates action strategies based on the current environment, ultimately completing the task.
@@ 5 Conclusion
In this paper, we have proposed a novel Robotic Perception and Planning framework (RoboMP) that consists of the Goal-Conditioned Multimodal Perceptor (GCMP) and the Retrieval-Augmented Multimodal Planner (RAMP). GCMP is introduced to capture multimodal environment information by incorporating a tailored MLLM. RAMP employs a coarse-to-fine retrieval-augmented approach to adapatively select the most-relevant policies as in-context demonstrations to enhance the generalization. Extensive experiments demonstrate that RoboMP outperforms the baselines by a large margin on both VIMABench and real-world tasks.
6 Impact Statements
Inspired by the great successes of multimodal large language models (MLLMs), we propose RoboMP that leverages both multimodal information in an environment and the general intelligence in MLLMs to enhance the perception and reasoning capabilities of embodied robots. To this end, this technology is expected to advance smart robots that can free human beings from some tedious work. There are many potential societal consequences of developing smart robots, none which we feel must be specifically highlighted here.
@@ Appendix D Planning Template
We construct a template to prompt multimodal generator to generate planning. The complete prompt is shown in Figure 7.

### v2

@@ 4.5 Ablation Study
Effects of the Coarse-to-fine Retriever. We conduct ablation experiments to valid the performance of RoboMP2 with different settings of the retriever-augmented module. In particular, we compare 4 different implementations: (1) w/o CR removes the coarse retriever, directly using the rewriting module to order all programs in the codebase; (2) w/o IR removes the rewriting module, remaining the coarse retriever and semantic reorder; (3) w/o SR removes the reorder module, unsorting the retrieved demonstrations; (4) w/o RAMP removes the whole retriever, and takes programs from the codebase randomly as prompt examples. As depicted in Table 4, without either of these components, the average performance drops significantly. The reason is that when the context is composed of irrelevant examples, it is difficult for the planner to generate execution plans successfully.
@@
| RoboMP2 | ✓ | ✓ | ✓ | 89.0 | 85.9 | 86.8 | 68.0 | 82.4 |
| RoboMP2 | ✗ | ✓ | ✓ | 14.4 | 16.0 | 19.6 | 8.0 | 14.5 |
| RoboMP2 | ✓ | ✗ | ✓ | 87.7 | 86.1 | 86.7 | 49.0 | 77.4 |
| RoboMP2 | ✓ | ✓ | ✗ | 81.9 | 80.6 | 82.7 | 65.0 | 77.5 |
| RoboMP2 | ✗ | ✗ | ✗ | 10.5 | 9.5 | 11.6 | 3.0 | 8.7 |
@@
To investigate the robustness of RoboMP2, we construct an extensive test set by enriching the original instructions using GPT4. Subsequently, we conduct experiments to evaluate the impact of this augmentation. As shown in Table 3, the results demonstrate the outstanding semantic understanding capability of our approach. In comparison to I2A, RoboMP2 is not confined to processing text instructions with fixed patterns. On the contrary, it possesses stronger generalization and robustness.
@@
Comparison of Multimodal Perceptors. To validate the capabilities of GCMP for intricate relationship references, we construct a perception test set. This set contains fundamental perception tasks that detect objects by names, and complex perception tasks that requires understanding complex referential expressions as illustrated in Section 3.1.1. Since the traditional visual perceptors used in the robotic manipulation, such as YOLO, cannot take the referential expressions to perceive objects, we compare our embodied perceptor GCMP with SAM+CLIP in I2A, and the general MLLM, i.e., Shikra (Chen et al., 2023). It can be seen from Table 7 that our GCMP outperforms the two models on both simple and complex referential perception by a large margin.
@@
Neither SAMP+CLIP or Shikra can work on the complex referential perception. The reason is that SAM+CLIP does not have a ability to understand the complex semantic reference and the general model Shikra struggles to generalize to complex embodied perception tasks.
@@ 4.6 Qualitative Results
We further present qualitative results in Figure 5, illustrating the planning and execution steps of RoboMP2 in comparison to other methods across three different kinds of tasks. It can be observed that only RoboMP2 successfully accomplishes all tasks. While all three methods perform well on VIMABench tasks, the end-to-end policy fails to generalize to real-world scenarios due to differences in the robot arm. Furthermore, SAM+CLIP schema is shown to unable to recognize referential objects with complex relationships. This leads I2A to only complete pick-and-place tasks for simple objects, resulting in failure in the second step of sim-to-real. In the generalization scenario, other methods are unable to complete the instruction. Only RoboMP2, leveraging its robust language understanding and multimodal perception capabilities, successfully identify objects with intricate relationships. It adaptively generates action strategies based on the current environment, ultimately completing the task.
@@ 5 Conclusion
In this paper, we have proposed a novel Robotic Perception and Planning framework (RoboMP2) that consists of the Goal-Conditioned Multimodal Perceptor (GCMP) and the Retrieval-Augmented Multimodal Planner (RAMP). GCMP is introduced to capture multimodal environment information by incorporating a tailored MLLM. RAMP employs a coarse-to-fine retrieval-augmented approach to adapatively select the most-relevant policies as in-context demonstrations to enhance the generalization. Extensive experiments demonstrate that RoboMP2 outperforms the baselines by a large margin on both VIMABench and real-world tasks.
Impact Statement
Inspired by the great successes of multimodal large language models (MLLMs), we propose RoboMP2 that leverages both multimodal information in an environment and the general intelligence in MLLMs to enhance the perception and reasoning capabilities of embodied robots. To this end, this technology is expected to advance smart robots that can free human beings from some tedious work. There are many potential societal consequences of developing smart robots, none which we feel must be specifically highlighted here.
Acknowledgements
We would like to thank the reviewers for their constructive comments. This work is supported by Shenzhen College Stability Support Plan (Grant No.GXWD20220817144428005). Additionally, it is also supported by China Postdoctoral Science Foundation (Grant No. 2023TQ0095 and 2023M740923) and partially supported by National Natural Science Foundation of China (Grant No.62306090).
@@ Appendix D Planning Template
We construct a template to prompt multimodal generator to generate planning. An example of complete prompt is shown in Figure 7. By combining the specific in-context sample, the length of the whole input tokens would be approximately 8k.